10-3(2-x)=x+8
solve for x

Answer by tara0066(31)

10-3(2-x)=x+8
10-(3)(2)-(3)(-x)=x+8 Distribute the 3 through the parenthesis.
10-6+3x=x+8
4+3x=x+8 Simplified
4+3x(-x)=x(-x)+8 Subtract x on both sides of the equation to get the variables together.
4+2x=8 Simplified
4(-4)+2x=8(-4) Subtract 4 from both sides to get x alone.
2x=4 Simplified
2x(/2)=4(/2) Divide both sides by 2
x=2